OpenStudy (anonymous):

When solving systems, if all of your numbers cancel out, is your answer 0 or 1?

OpenStudy (skullpatrol):

It depends on what you are left with.

OpenStudy (anonymous):

Example: 4x-y=1 and -4x+y=-1

OpenStudy (anonymous):

I got 4x-2=1 and -4x+2=-2 so all of the coefficients would cancel out so would x=0?

OpenStudy (skullpatrol):

No.

OpenStudy (anonymous):

So x would equal 1 then?

OpenStudy (skullpatrol):

If you are left with a true statement such as 1 = 1 then your solution set is the set of all real numbers, if you are left with a false statement such as 2 = 1 then your solution set is the empty set.

OpenStudy (skullpatrol):

In your example: 4x-y=1 and -4x+y=-1 are the equivalent equations since the second is -1 times the first.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Its just asking you to solve for x and y

OpenStudy (skullpatrol):

Look at the equations, Example: 4x-y=1 and -4x+y=-1

OpenStudy (anonymous):

Wait i think i get it now

OpenStudy (anonymous):

thanks!

OpenStudy (skullpatrol):

$$ 4x-y=1 $$ $$-1(4x-y)=-1(1)$$ $$-4x+y=-1$$

OpenStudy (anonymous):

ok thanks!

OpenStudy (skullpatrol):

Thanks for trying.